what is 1/3 of $58900.= 1*58900/3=19633.33 $
Found 2 solutions by checkley77, mangeshpjoshi:
Answer by checkley77(12844)

what is 1/3 of $58900.
1/3*58,900=58,900/3=19,633 1/3 ans.

Answer by mangeshpjoshi(15)

1*58900/3=19633.33$ .33 is recurring factor ,
beside this is correct .
faithfully
mangesh joshi